these two questions im having trouble on. Provides the tables for both
M 6.2.75 се Use a standard normal distribution table to obtain the Z-score that has an area of 0.975 to its right. Click here normal distribution page 1 Areas under the standard normal curve Second decimal place in 2 0.06 0.05 0.04 0.03 0.09 0.08 0,07ethods vork: Chapter 6 pt 11 of 20 (14 complete) i normal distribution page 2 Areas under the standard normal curve Second de



M 6.2.75 CC Use a standard normal distribution table to obtain the Z-score that has an area of 0.975 to its right. Click herenormal distribution page 1 Second decimal place in 2 Areas under the standard normal curve 0.09 0.08 0,07 0.06 0.05 0.04 0.03Second decimal place in z TABLE II (cont.) Areas under the standard normal curve 0.00 0.01 0.02 Z 0.03 0.04 0.05 0.06 0.07 0.
Answer #1

6.2.75

We have to obtain z score which has area 0.975 to its right.

We are given that,

\small P(Z>a)=0.975    where a is the required z score.

\small or, \ 1-P(Z\le a)=0.975

\small or, \ P(Z\le a)=0.025

\small or, \ \Phi(a)=0.025=\Phi(-1.96) \ \ (Value \ from \ z-table)

Therefore, value of a is -1.96.     (Answer)